OpenStudy (anonymous):

A deck of cards contains 6 black cards numbered 5 to 10 and 6 red cards numbered 1 to 6. How many 5-card hands having 3 black cards and 2 red cards can be dealt?

OpenStudy (markchernioglo):

What do you think?

OpenStudy (anonymous):

I think it is 65 but not positive

OpenStudy (kropot72):

The number of combinations of the 6 black cards, taken 3 at a time, is given by \[\large 6C3=\frac{6!}{3!3!}\] The number of combinations of the 6 red cards, taken 2 at a time, is given by \[\large 6C2=\frac{6!}{2!4!}\] Each combination of black cards can be taken with each combination of red cards. Therefore the total number of combinations is given by \[\large 6C3 \times 6C2=you\ can\ calculate\]

OpenStudy (anonymous):

I got 300, thanks!

OpenStudy (kropot72):

You're welcome :)